From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 18 2003 - 08:15:07 EST
------------------------------------------------------------ Date: Tue Mar 18 08:15:06 2003, Shift: 00:00-08:00 Supervisor: Michael Murray, with: Svein Ledal Log book#14 pages 67 - 67 Shift Summary: ------------------------------------------------------------ 1) Machine off till about 2pm. Brahms got beam very late and then we found that we had a HV power surply in the zone that could not be restarted with an access. MCR said they would dump at 3.45. Beam back again at 6am. 2) We took just a small amount of data at full field and 4 degrees in the FS. Then moved to 8 degrees and took 120K trig 6 at 1/4B field. 8485 Run 8486 is 1/8B field, currently it has 160K of trigger 6. 3) Apart from HV surply no problems with detectors. Dana wants to fix 1 slat with his trigger counters. ***** Run 8478 was actually 1/4B at 8 degrees ***** Runs 8485,8486 were actually 1/8th B at 8 degrees ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Tue Mar 18 2003 - 08:21:44 EST